The probability that an open heart operation is successful is 0.84. what is the probability that in two random selected open heart operations, at least one will be successful?
Answer by Fombitz(32388)

4 possible outcomes.
1: Success, 2:Success, P=0.84(0.84)=0.7056
1:Success, 2: Failure, P=0.84(0.16)=0.1344
1:Failure, 2:Success, P=0.16(0.84)=0.1344
1:Failure, 2:Failure, P=0.16(0.16)=0.0256
At least one success would be the total of the first three,
P(at least 1 success)=0.7056+0.1344+0.1344=0.9744